### v1

@@ 5 CONCLUSION
Limitations and Future Works. Previous work (Yosinski et al., 2014; Raghu et al., 2017) has suggested that different layers within a network exhibit varying training dynamics, which also relates to model memory. Shallow layers tend to learn general knowledge, while deeper layers tend to learn task-specific knowledge, with memory primarily occurring in deeper layers (Stephenson et al., ). We have not yet explored this aspect. Designing or exploring the memory mechanisms and memory terms for structurally different models or different layers within a model is one of our future research directions.
@@ B.1 Image Classification Dataset
CIFAR10. The CIFAR-10 dataset (Krizhevsky et al., 2010) consists of 60,000 32x32 color images, divided into 50,000 training images and 10,000 test images. The dataset is divided into 10 classes, including: airplanes, cars, cats, etc. Each class contains 6,000 images.
@@ B.2 Natural Language Datasets
GLUE. The General Language Understanding Evaluation (GLUE) benchmark (Wang, 2018) consists of nine natural language understanding (NLU) tasks, all in English. The GLUE tasks include single-sentence tasks like CoLA and SST-2, similarity and paraphrase tasks such as MRPC, STS-B, and QQP, as well as natural language inference tasks including MNLI, QNLI, RTE, and WNLI. Well-known models like BERT (Devlin, 2018), XLNet (Yang, 2019), RoBERTa (Delobelle et al., 2020), ERINE, T5, and others are tested on this benchmark. Specific details of each dataset are reported in Table (B.1).
@@ C.1 Supervised Learning
In supervised learning, we conducted experiments on the CIFAR-10/100 and ImageNet-1K datasets. In CIFAR, we used the ResNet18 and ResNet34 models, along with the LARS optimizer (You et al., 2017) and OneCycle learning rate schedule (Smith and Topin, 2017). For ImageNet-1K, we employed the ResNet50 model, and the learning rate was multiplied by 0.1 every 30 epochs. Our other hyperparameters are reported in Table B.2 and B.3.
200 0.2 128 0.9 5e-4 LARS(You et al., 2017) 5.2 OneCycle(Smith and Topin, 2017)
Epoch lr batch size momentum weight decay optimizer
90 0.1 256 0.9 1e-4 SGD
@@
Table B.4: The hyperparameters we used in the GLUE experiments.
Epoch lr batch size optimizer lr scheduler
10 2e-5 32 Adam(Kingma and Ba, 2014) linear
@@ C.2 Self-Supervised Learning
    Table B.5: The hyperparameters we used in SimCLR.
    Pre-training
    Epoch lr batch size optimizer weight decay
    100 1e-3 256 Adam(Kingma and Ba, 2014) 1e-6
    Fine-tuning
    Epoch lr batch size optimizer weight decay
    100 1e-3 32 Adam(Kingma and Ba, 2014) 1e-6
[…]

### v2

@@ 5 Conclusion
[…]
Declarations
  • Funding. None
  • Conflict of interest/Competing interests (check journal-specific guidelines for which heading to use). The authors reported no potential conflict of interest.
  • Ethics approval and consent to participate. Not applicable.
  • Consent for publication. All authors consent to the publication of this work.
  • Data availability. The data used in this work are all open-source datasets
  • Materials availability Not applicable.
  • Code availability. The code was shown in https://github.com/xiaojinying/EMP
  • Author contribution. JinYing Xiao conceived the research idea and developed the study design. Ping Li wrote the main manuscript text, developed the methodology, and contributed to software implementation. Jie Nie validated the results and contributed to software development. Zhe Tang prepared the visualizations. All authors reviewed and edited the manuscript.
@@ B.1 Image Classification Dataset
CIFAR10. The CIFAR-10 dataset [ref54] consists of 60,000 32x32 color images, divided into 50,000 training images and 10,000 test images. The dataset is divided into 10 classes, including: airplanes, cars, cats, etc. Each class contains 6,000 images.
@@ B.2 Natural Language Datasets
GLUE. The General Language Understanding Evaluation (GLUE) benchmark [ref35] consists of nine natural language understanding (NLU) tasks, all in English. The GLUE tasks include single-sentence tasks like CoLA and SST-2, similarity and paraphrase tasks such as MRPC, STS-B, and QQP, as well as natural language inference tasks including MNLI, QNLI, RTE, and WNLI. Well-known models like BERT [ref24], XLNet [ref52], RoBERTa [ref53], ERINE, T5, and others are tested on this benchmark. Specific details of each dataset are reported in Table (B.1).
@@
200 0.2 128 0.9 5e-4 LARS([ref55]) 5.2 OneCycle([ref56])
Epoch lr batch size momentum weight decay optimizer
90 0.1 256 0.9 1e-4 SGD
Table B.4: The hyperparameters we used in the GLUE experiments.
Epoch lr batch size optimizer lr scheduler
10 2e-5 32 Adam([ref57]) linear
Table B.5: The hyperparameters we used in SimCLR.
Pre-training
Epoch lr batch size optimizer weight decay
100 1e-3 256 Adam([ref57]) 1e-6
Fine-tuning
Epoch lr batch size optimizer weight decay
100 1e-3 32 Adam([ref57]) 1e-6
@@ C.1 Supervised Learning
In supervised learning, we conducted experiments on the CIFAR-10/100 and ImageNet-1K datasets. In CIFAR, we used the ResNet18 and ResNet34 models, along with the LARS optimizer [ref55] and OneCycle learning rate schedule [ref56]. For ImageNet-1K, we employed the ResNet50 model, and the learning rate was multiplied by 0.1 every 30 epochs. Our other hyperparameters are reported in Table B.2 and B.3.